OpenStudy (anonymous):

Ellis Klein borrows $2,000 at an annual interest rate of 7.2% for two years. Calculate his monthly payments

OpenStudy (anwaarullah):

see answer in attached file,,,,,2000*7.20%=144 for two year so 144*2=288 principal with interest in two years......288+2000=2,200 for one month divide by 12......2,288/12= 190.67 per month with interest